Find [theta] for 0° <= 0<360°
sin[theta]=-0.8412

Answer by stanbon(75887)

Find [theta] for 0°<= theta <360°
sin[theta]=-0.8412
-----------------
sin^-1(-0.8412) = -57.26
---
theta = 360-57.26 = 302.74 degrees
OR
theta = 57.26+180 = 237.26 degrees
======================================
Cheers,
Stan H.